OpenStudy (anonymous):

Determine algebraically whether the function is even, odd, or neither even nor odd. f(x)=2/x2 (the x2 represents an exponent) Neither Even Odd

OpenStudy (anonymous):

?\[f(x)=\frac{2}{x^2}\]?

OpenStudy (perl):

is that the function?

OpenStudy (anonymous):

satellite has it correct

OpenStudy (perl):

the definition of even function is : if f(-x) = f(x) for any x, then f is even function

OpenStudy (perl):

f(-x) = 2 / (-x)^2 = 2 / x^2 = f(x) so f is even

OpenStudy (anonymous):

thanks :D